jordan opened a one-gallon container of milk and poured one pint of milk into his glass. what is the fractional part of one gallon left in the container ?
a 1/8 b 1/2 c 3/4 d 7/8

Answer by mananth(16949)

1 pint = 1/8 gallon
1-1/8 = 7 /8 gallon remains